# Learning Safe Strategies for Value Maximizing Buyers in Uniform Price Auctions

### v2

@@ D.2.3.  Computing $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$
D.3.  Tight Lower Bound for Theorem 5.4
@@
Table 3: Example of bid history, $\mathcal{H}_{-}$, achieving 4-approximation asymptotically for $m^{\prime}=4$ (and $m=1$). Each round in the same partition has identical competing bid profile. Total number of units in each auction is $K=N^{3}$.
@@
D.3.2.  Computing $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$
We make the following claim about the optimal $m^{\prime}$-uniform safe bidding strategy for $\mathcal{H}_{-}$.
@@ D.3.2.  Computing $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$
D.3.3.  Computing $\mathbf{b}_{1}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$
In this section, we compute the optimal 1-uniform safe bidding strategy for the bid history, $\mathcal{H}_{-}$.
@@ D.3.3.  Computing $\mathbf{b}_{1}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$
Substituting values from Lemma D.3, Lemma D.4 and Eq. 48
$\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{1}{m}\cdot\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ $=\frac{V^{\mathsf{SAFE}}_{m^{\prime}}\left(\mathcal{H}_{-}\right)}{mV^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}$
@@
D.4.  Tight Lower Bound for Theorem 5.5
@@
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{1}{m}\cdot\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ $=\frac{N^{2m^{\prime}-1}+(2m^{\prime}-1)(N^{2m^{\prime}-1}-N^{2m^{\prime}-2})v}{mN^{2m^{\prime}-1}}$

### v4

@@ D.1.3. Computing $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$
D.2. Tight Lower Bound for Theorem 5.4
@@
Table 4: Bid history attaining tight lower bound for $m^{\prime}=4$ (and $m=1$). Each round in the same partition has identical competing bid profile. Total number of units in each auction is $K=N^{3}$.
@@
D.2.2. Computing $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$
@@
D.2.3. Computing $\mathbf{b}_{1}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$
@@
Substituting values from Lemma D.3, Lemma D.4
[…]
$=\frac{V^{\mathsf{SAFE}}_{m^{\prime}}\left(\mathcal{H}_{-}\right)}{mV^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}$
@@
D.3. Tight Lower Bound for Theorem 5.5
@@
[…]
$=\frac{N^{2m^{\prime}-1}+(2m^{\prime}-1)(N^{2m^{\prime}-1}-N^{2m^{\prime}-2})v}{mN^{2m^{\prime}-1}}$